OpenStudy (anonymous):

The velocity of a particle moving along the x-axis is: v(t)= -(t-1)(^2) + 2... for t between 0 and 6. if at t=0 the particle is at the origin, find the location of the particle at t=3

OpenStudy (anonymous):

\[\textbf v(t) = \frac{\mathrm{d}\textbf x(t)}{\mathrm{d}t}\] \[\int_{0}^{t≤6s} \textbf v(t) \mathrm{d}t=\int_{x_0=0}^{x}\mathrm{d}x\] \[\begin{array}{l} \textbf x(3) & =\int_{0}^{3} \big(-(t-1)^2+2 \big) \ \mathrm{d}t \\ & = \int_{0}^{3}(1+2t-t^2) \mathrm{d}t \end{array}\] Then integrate Remember for the sake of completeness that the units on velocity are m/s, and the units on position are meters, so technically the units on each term of the integration expression is \[\int_{0}^{t}(1m/s)+(2m/s^2)t-(1m/s^3)t^2) \mathrm{d}t\] This doesn't change the numeric answer, but is good to think about.

OpenStudy (anonymous):

Make sense? ^_^

OpenStudy (anonymous):

yes thank you!

OpenStudy (anonymous):

^_^